Two cars A and B were at town P. They wanted to travel to town Q. Car A started 15 minutes earlier than car B and maintained a constant speed of 20 m/s whereas car B started from rest and maintained a constant acceleration. Car A took 30 minutes to reach the destination. If car A and car B reached the destination simultaneously, then velocity with which car B reached is 


1
Expert's answer
2021-08-02T09:59:32-0400

Given:

"t_A=30\:\rm min=1800\: s"

"t_B=15\:\rm min=900\: s"

"v_A=20\:\rm m/s=const"

"v_{Bi}=0"

The distance between Q and P

"d=v_A*t_A=1800*20=36000\:\rm m"

From the other side

"d=\frac{v_{Bi}+v_{Bf}}{2}*t_B"

Hence, the final velocity of car B

"v_{Bf}=\frac{2d}{t_B}=\frac{2*36000}{900}=40\:\rm m/s"